A new temperature scale is called Z. On that scale, the boiling point of water is 150 degrees Z and the freezing point is −10.0 degrees Z . Find the corresponding 70 degrees Z in degrees Celsius. they said that the answer is 50 but i don't know the solution
Answer by macston(5194)

.
From the freezing point of water to the boiling point is 100 centigrade degrees and 160 Z degrees. 100 C degrees/160 Z degrees=5/8 C degree per Z degree
The Z scale starts 10 degrees lower, so to zero this we add 10.
To convert from Z to C:
C=5/8(Z+10 degrees)
To convert from C to Z:
Z=((8/5)C)-10
Notice we add 10 to Z first, going from Z to C we subtract after.
.
C=5/8(Z+10)=5/8(70+10)=5/8(80)=50 degrees C
.
CHECK:
Z=((8/5)C)-10
70=((8/5)50)-10
70=80-10
70 =70